# Search for spatial coincidences between galaxy mergers and Fermi-LAT 4FGL-DR4 sources

### v1

@@ Abstract
We present a systematic search for spatial association between a high-confidence sample of 3,166 morphologically selected galaxy mergers detected through SDSS imaging and the 4FGL-DR4 catalog of gamma-ray sources detected by the Fermi Large Area Telescope (LAT). Using a conservative 4 positional uncertainty threshold and a Poisson-based statistical framework, we identify 21 statistically significant associations with match probabilities . Among these are known classes of gamma-ray emitters such as flat-spectrum radio quasars (FSRQs), BL Lacertae objects (BLL), and radio galaxies bolstering the hypothesis that merger-driven processes can fuel high-energy activity. Intriguingly, five of the associated sources remain unclassified in 4FGL-DR4, hinting at a possible link between galaxy mergers and a hitherto unrecognized population of gamma-ray sources. The dominance of AGN-like matches supports the scenario in which mergers trigger accretion onto central supermassive black holes, initiating AGN activity observable at gamma-ray energies. Moreover, the recurrent presence of unassociated sources among secure matches underscores the potential of merger catalogs as physically motivated priors in gamma-ray source identification efforts. This work constitutes the first dedicated effort to explore associations between gamma-ray sources and a large, morphologically selected sample of galaxy mergers, opening new avenues for understanding the role of interactions in high-energy astrophysics. We additionally examined the sample of 70 galaxy pairs from the Canadian Network for Observational Cosmology Field Galaxy Redshift Survey and found no statistically significant matches, with within the 4 positional uncertainty threshold.
I Introduction
@@
The Fermi Large Area Telescope (LAT) has been a cornerstone in the exploration of the gamma-ray universe at GeV energies since its launch in 2008 Atwood et al. (2009) 111https://fermi.gsfc.nasa.gov/science/instruments/lat.html. It has provided an unprecedented view of the high-energy sky, enabling the detection and characterization of thousands of gamma-ray sources. The most recent data release, the Fourth Fermi-LAT Source Catalog Data Release 4 (4FGL-DR4) provides a comprehensive catalog of 7,195 gamma-ray sources observed over 14 years from 50 MeV to 1 TeV Ballet et al. (2023) (see also the related 4FGL-DR3 catalog Abdollahi et al. (2022)). These sources include a variety of astrophysical objects, such as blazars, radio galaxies, pulsars. However, a large number of Fermi-LAT sources () remain unidentified and no counterparts at lower energies have been identified. Numerous studies have therefore searched for spatial associations of Fermi 4FGL-DR4 sources with multi-wavelength source catalogs, including X-ray catalogs from SRG/eROSITA Mayer and Becker (2024) and SWIFT Ulgiati et al. (2025), radio source catalogs Bruzewski et al. (2023), SPT point sources at millimeter wavelengths Zhang et al. (2022), magnetars Ramakrishnan and Desai (2024), IceCube neutrino events Negro et al. (2023), MeerKAT sources Himes et al. (2025) etc.
@@
In this work, we search for statistically significant association between Fermi 4FGL-DR4 catalog and galaxy mergers. Galaxy mergers are pivotal events in the evolution of galaxies, driving significant morphological and dynamical changes that influence star formation, black hole growth, and active galactic nuclei (AGN) activity Hopkins et al. (2006); Kaviraj (2025). These processes can produce energetic phenomena at gamma-ray energies, which can be detected by LAT. Galaxy mergers are thought to trigger intense star formation and fuel supermassive black holes, potentially leading to relativistic jets or other mechanisms responsible for gamma-ray production. Understanding the potential connection between galaxy mergers and gamma-ray sources is crucial for unraveling the physical processes governing these extreme astrophysical environments. A systematic search for associations between a comprehensive, high-confidence catalog of galaxy mergers and the latest Fermi-LAT data release has not yet been done, to the best of our knowledge.
@@
This work is heavily inspired by Bouri et al. (2025), which searched for a spatial association between IceCube-detected neutrinos and sources from six different galaxy merger catalogs, since neutrinos could be produced through shock acceleration from galaxy mergers, as discussed in various works Lisenfeld and Völk (2010); Kashiyama and Mészáros (2014); Yuan et al. (2018). Along with the associated neutrinos, one would also expect high-energy gamma-ray radiation, which could be detected by Fermi-LAT. We briefly review some of the gamma-ray production mechanisms discussed in literature. The first estimate of gamma-ray flux from galaxy merger candidates was done in Lisenfeld and Völk (2010). This work considered two specific merger systems, viz. UGC 12914/5 and UGC 813/6 Lisenfeld and Völk (2010). These candidates are located at distances of 60 Mpc and 69 Mpc, respectively. The hadronic gamma-ray flux from decays resulting from energetic proton collisions with gas nuclei in the shocked region has been estimated in this fork to be , with a spectral index of 1.1 normalized at 1 TeV Lisenfeld and Völk (2010). Although this flux is too small to be detected by Fermi-LAT, for a galaxy merger happening at distances between 2.6 and 3.9 Mpc, this flux could be up to 700 times higher, thereby enhancing the detection probability of mergers in current generation gamma-ray telescopes such as Fermi-LAT. Then, Kashiyama and Mészáros (2014) investigated the shock acceleration of particles in galaxy mergers with stellar masses of and showed that cosmic rays could be accelerated upto 0.1-1 EeV with a spectral index of 2.0. It was shown that for a cosmic ray acceleration efficiency of 10%, one could detect gamma-ray radiation from about 10 nearby merging galaxies with Cherenkov Telescope Array Kashiyama and Mészáros (2014). In addition to UGC 12914/5 and UGC 813/6, another galaxy merger candidate, viz. VV 114 located at a distance of Mpc Vorontsov-Velyaminov et al. (2001) was identified as a potential gamma ray source. Subsequently, Yuan et al. (2018) extended this analysis and calculated the cumulative diffuse gamma-ray background by incorporating mergers of galaxy cluster and group candidates, and also incorporated the redshift evolution of average galactic radius, shock velocity and gas content inside haloes and galactic magnetic fields. This model is also consistent with the non-blazar contribution to extragalactic gamma-ray background Yuan et al. (2018).
The empirical study of galaxy mergers and correctly identifying galaxies that have been merging has traditionally been quite challenging (See Kaviraj (2025) for a recent review). Recent advances in catalogs from large-scale optical surveys, such as the Sloan Digital Sky Survey Data Release 16 (SDSS DR16) Ahumada et al. (2020), have however enabled the identification of large samples of galaxy mergers through morphological and spectroscopic signatures. Using machine learning techniques and high-confidence classification thresholds, Ackermann et al. (2018) constructed a galaxy merger catalog. This catalog contains a high-confidence sample of 3,166 galaxy mergers derived from a catalog of 328,151 candidates in the redshift range Ackermann et al. (2018). The second catalog we considered for our analysis is the J/AJ/130/2043 catalog, from the Canadian Network for Observational Cosmology field galaxy redshift survey, contains 70 close galaxy pairs across four sky regions covering 1.5 square degrees, with redshifts in the range Patton et al. (2005). We use both these aformentioned catalogs for spatial coincidence with Fermi-LAT point sources. These catalogs were also used for spatial coincidence analysis with IceCube neutrinos in Bouri et al. (2025). We note that other major merger catalogs Hwang and Chang (2009); Gimeno et al. (2004) used in Bouri et al. (2025) lack essential photometric or distance information required for flux-based probability estimates. Furthermore, several prominent merger systems discussed in the literature as promising gamma-ray sources, including UGC 12914/5, UGC 813/6, VV 114, NGC 660, and NGC 3256, are not present in the galaxy merger catalogs used in this study Ackermann et al. (2018). We also searched for spatial associations of these systems with sources in the 4FGL-DR4 catalog within a radius but found no matches. A targeted analysis of such individual systems will be conducted in a follow-up study.
II Data Analysis
@@
We analyze the aformentioned galaxy merger catalog that contains 328,151 candidate mergers Ackermann et al. (2018). Each galaxy image in the catalog was processed by a convolutional neural network (CNN) classifier, which outputs a merger probability score , representing the likelihood that the system is undergoing a merger (with indicating a non-interacting galaxy and indicating a merger). The classifier was trained using 4-fold cross-validation, producing four independent estimates of for each galaxy Ackermann et al. (2018). To isolate high-confidence systems, we applied a strict threshold of across all four classifier outputs, retaining only those galaxies that satisfied this criterion in every fold. This conservative approach yielded a subset of 3,711 robust merger candidates. This stricter criterion minimizes the inclusion of ambiguous cases and emphasizes classifier agreement, albeit at the cost of reducing the final sample size. These selected objects were then cross-matched with the SDSS DR16 catalog Ahumada et al. (2020) using TOPCAT version 4.10-3 Taylor (2005), employing a matching radius of 0.1 arcseconds. The resulting cross-match identified 3,185 mergers with SDSS counterparts. Among these matched sources, 1,147 have measured spectroscopic redshifts and 3,166 sources possess SDSS -band apparent magnitudes. To characterize the spatial extent of the merger sample, we estimated the total angular sky area over which the 3,166 galaxy mergers are distributed. This effective footprint was computed by dividing the sky into finer bins in right ascension and declination and summing up the area of all the spatial bins that contained at least one merger. The contribution of each bin was corrected for spherical projection by weighting with cos(), where is the central declination of the bin. This produced an estimated sky coverage of approximately , which was used for surface density estimates and association probability calculations. Figure 1 shows the sky distribution in equatorial coordinates using an Aitoff projection.
@@
For the gamma-ray dataset, we utilized 4FGL-DR4 catalog, which contains 7,195 gamma-ray sources detected between 50 MeV and 1 TeV over 14 years of observations Ballet et al. (2023).
@@
To identify possible associations, we searched for candidate galaxy mergers located within the 4 positional uncertainty ellipse of each 4FGL-DR4 source. This conservative threshold, similar to that used in Zhang et al. (2022), accounts for both Fermi-LAT’s positional uncertainties and uncertainties in the external catalogs. When multiple optical sources were found within a single gamma-ray error ellipse, we assessed the probability of random association using a Poisson-based framework as described below Browne and Cohen (1978); Downes et al. (1986); Biggs et al. (2011); Zhang et al. (2022).
@@
The expected number of chance associations ) is given by Browne and Cohen (1978); Downes et al. (1986); Biggs et al. (2011); Zhang et al. (2022):
@@
where and are the semi-major and semi-minor axes of the Fermi positional ellipse, and is the surface density of mergers with flux greater than that of the SDSS counterpart having flux equal to . To evaluate , we used the SDSS-derived fluxes computed from the -band magnitudes of the merger candidates. For each candidate to be tested, we determined the number of mergers in the full sample with -band fluxes exceeding the corresponding candidate flux value. The details of the calculation of -band flux from the estimated magnitude can be found in Appendix A. Dividing this count by the total sky area covered by the catalog yielded a flux-dependent surface density. This value of was then used in Eq. 1 to calculate the expected number of chance associations (), allowing us to assess the probability of random coincidence.
@@
A small value indicates a low probability of random association, which implies a more reliable counterpart match. We adopt a threshold of to define a “secure” association, with a secondary “relaxed” threshold of also examined, following the criteria in Zhang et al. (2022). Given the rarity and low surface density of galaxy mergers, we expect the false match rate to remain low even with a relaxed threshold Downes et al. (1986); Biggs et al. (2011).
III Results

### v2

@@ Abstract
We present a systematic search for spatial association between a high confidence sample of 3,166 morphologically selected galaxy mergers drawn from the initial convolutional neural network (CNN) catalog of 328,151 candidate mergers (Ackermann et al., 2018) and refined through SDSS cross-matching and the Fermi Large Area Telescope Fourth Source Catalog Data Release 4 (4FGL-DR4) catalog of gamma-ray sources detected by the Fermi Large Area Telescope (LAT). Using a conservative 4 positional uncertainty threshold and a Poisson based statistical framework, we identify 21 statistically significant associations with match probabilities . Among these are known classes of gamma-ray emitters such as Flat-Spectrum Radio Quasars (FSRQs), BL Lacertae objects (BLL), and radio galaxies bolstering the hypothesis that merger driven processes can fuel high energy activity. Intriguingly, five of the associated sources remain unclassified in 4FGL-DR4, hinting at a possible link between galaxy mergers and a hitherto unrecognized population of gamma-ray sources. The dominance of AGN like matches supports the scenario in which mergers trigger accretion onto central supermassive black holes, initiating AGN activity observable at gamma-ray energies. Moreover, the recurrent presence of unassociated sources among secure matches underscores the potential of merger catalogs as physically motivated priors in gamma-ray source identification efforts. This work constitutes the first dedicated effort to explore associations between gamma-ray sources and a large, morphologically selected sample of galaxy mergers (Ackermann et al., 2018) catalog, opening new avenues for understanding the role of galaxy interactions in high-energy astrophysics. We additionally examined the sample of 70 galaxy pairs from the Canadian Network for Observational Cosmology Field Galaxy Redshift Survey and found no statistically significant matches, with within the 4 positional uncertainty threshold.
keywords:
Gamma-ray astronomy; Galaxy mergers1 Introduction
@@
The Fermi Large Area Telescope (LAT) has been a cornerstone in the exploration of the gamma-ray universe at GeV energies since its launch in 2008 (Atwood et al., 2009) 111https://fermi.gsfc.nasa.gov/science/instruments/lat.html. It has provided an unprecedented view of the high energy sky, enabling the detection and characterization of thousands of gamma-ray sources. The most recent data release, the Fourth Fermi-LAT Source Catalog Data Release 4 (4FGL-DR4) provides a comprehensive catalog of 7,195 gamma-ray sources observed over 14 years from 50 MeV to 1 TeV (Ballet et al., 2023) (see also the related 4FGL-DR3 catalog Abdollahi et al. 2022). These sources include a plethora of astrophysical objects, such as blazars, radio galaxies, pulsars, etc. However, a large number of Fermi-LAT sources () remain unidentified and no counterparts at lower energies have been identified. Numerous studies have therefore searched for spatial associations of Fermi 4FGL-DR4 sources with multi-wavelength source catalogs, including X-ray catalogs from SRG/eROSITA (Mayer and Becker, 2024) and SWIFT (Ulgiati et al., 2025), radio source catalogs (Bruzewski et al., 2023), South Pole Telescope (SPT) point sources at millimeter wavelengths (Zhang et al., 2022), magnetars (Ramakrishnan and Desai, 2025), IceCube neutrino events (Negro et al., 2023), MeerKAT sources (Himes et al., 2025) etc.
@@
In this work, we search for statistically significant association between Fermi 4FGL-DR4 catalog and galaxy mergers. Galaxy mergers are pivotal events in the evolution of galaxies, driving significant morphological and dynamical changes that influence star formation, black hole growth, and active galactic nuclei (AGN) activity (Hopkins et al., 2006; Kaviraj, 2025). These processes can produce energetic phenomena at gamma-ray energies, which can be detected by LAT. An important observational precedent for this scenario was reported by Paliya et al. (2020), who presented the first definitive case of a galaxy merger (TXS 2116077) hosting a young, relativistic gamma-ray jet in a narrow line Seyfert 1 galaxy. This provides strong empirical support for the idea that merger triggered AGN activity can give rise to high energy gamma-ray emission (Paliya et al., 2020; Ramos Almeida et al., 2013; Chiaberge et al., 2015).
Galaxy mergers are also thought to trigger intense star formation and fuel supermassive black holes, potentially leading to relativistic jets or other mechanisms responsible for gamma-ray production (Ellison et al., 2008, 2011; Scudder et al., 2012; Satyapal et al., 2014; Gabor et al., 2016; Kashiyama and Mészáros, 2014). Understanding the potential connection between galaxy mergers and gamma-ray sources is crucial for unraveling the physical processes governing these extreme astrophysical environments. A systematic search for associations between a comprehensive, high confidence catalog of galaxy mergers and the latest Fermi-LAT data release has not yet been done, to the best of our knowledge.
@@
This work is heavily inspired by Bouri et al. (2025), which searched for a spatial association between IceCube-detected neutrinos and sources from six different galaxy merger catalogs, since neutrinos could be produced through shock acceleration from galaxy mergers, as discussed in various works (Lisenfeld and Völk, 2010; Kashiyama and Mészáros, 2014; Yuan et al., 2018). Along with the associated neutrinos, one would also expect high energy gamma-ray radiation, which could be detected by Fermi-LAT. We briefly review some of the gamma-ray production mechanisms discussed in literature. The first estimate of gamma-ray flux from galaxy merger candidates was done in Lisenfeld and Völk (2010). This work considered two specific merger systems, viz. UGC 12914/5 and UGC 813/6 (Lisenfeld and Völk, 2010). These candidates are located at distances of 60 Mpc and 69 Mpc, respectively. The hadronic gamma-ray flux from decays resulting from energetic proton collisions with gas nuclei in the shocked region has been estimated in this fork to be , with a spectral index of 1.1 normalized at 1 TeV (Lisenfeld and Völk, 2010). Although this flux is too small to be detected by Fermi-LAT, for a galaxy merger happening at distances between 2.6 and 3.9 Mpc, this flux could be up to 700 times higher, thereby enhancing the detection probability of mergers in current generation gamma-ray telescopes such as Fermi-LAT.
Then, Kashiyama and Mészáros (2014) investigated the shock acceleration of particles in galaxy mergers with stellar masses of and showed that cosmic rays could be accelerated up to 0.1-1 EeV with a spectral index of 2.0. It was shown that for a cosmic ray acceleration efficiency of 10%, one could detect gamma-ray radiation from about 10 nearby merging galaxies with Cherenkov Telescope Array (Kashiyama and Mészáros, 2014). In addition to UGC 12914/5 and UGC 813/6, another galaxy merger candidate, viz. VV 114 located at a distance of Mpc (Vorontsov-Velyaminov et al., 2001) was identified as a potential gamma ray source. Subsequently, Yuan et al. (2018) extended this analysis and calculated the cumulative diffuse gamma-ray background by incorporating mergers of galaxy cluster and group candidates, and also incorporated the redshift evolution of average galactic radius, shock velocity, and gas content inside haloes and galactic magnetic fields. This model is also consistent with the non-blazar contribution to extragalactic gamma-ray background (Yuan et al., 2018). The empirical study of galaxy mergers and correctly identifying galaxies that have been merging has traditionally been quite challenging (See Kaviraj 2025 for a recent review). Recent advances in catalogs from large-scale optical surveys, such as the Sloan Digital Sky Survey Data Release 16 (SDSS DR16) (Ahumada et al., 2020), have however enabled the identification of large samples of galaxy mergers through morphological and spectroscopic signatures.
The first catalog analyzed in this work is based on the convolutional neural network (CNN) catalog of 328,151 candidate mergers identified by Ackermann et al. (2018) from which we filtered 3,166 galaxy mergers for our analysis; further details of which are provided in Section 2. The second catalog we considered for our analysis is the J/AJ/130/2043 catalog, from the Canadian Network for Observational Cosmology field galaxy redshift survey, contains 70 close galaxy pairs across four sky regions covering 1.5 square degrees, with redshifts in the range (Patton et al., 2005). We use both these aformentioned catalogs for spatial coincidence with Fermi-LAT point sources. These catalogs were also used for spatial coincidence analysis with IceCube neutrinos in Bouri et al. (2025).
We note that other major merger catalogs (Hwang and Chang, 2009; Gimeno et al., 2004) used in Bouri et al. (2025) lack essential photometric or distance information required for flux-based probability estimates. Several well studied galaxy merger systems, such as UGC 12914/5, UGC 813/6, and VV 114, have been identified in prior literature as potential gamma ray sources due to their enhanced star formation and AGN activity triggered by gravitational interactions (Lisenfeld and Völk, 2010; Kashiyama and Mészáros, 2014). Similarly, systems like NGC 660 and NGC 3256 are noted for their intense starburst and AGN driven processes, which could contribute to gamma-ray emission (Ackermann et al., 2018). However, these systems are not included in the galaxy merger catalogs used in this study (Ackermann et al., 2018). We searched for spatial associations between these systems and sources in the 4FGL-DR4 catalog within 1 arcminute radius, based on the positional accuracy of the catalog, but found no matches. A targeted analysis of these individual systems is being conducted in a follow-up study to further explore their gamma-ray emission potential.
While earlier studies have explored potential links between galaxy mergers and high-energy neutrino emission (Bouri et al., 2025), no systematic effort has yet been made to investigate such connections in the gamma-ray regime. Our work represents the first dedicated search for associations between gamma-ray sources and a large, morphologically selected sample of galaxy mergers (Ackermann et al., 2018), thereby opening new avenues for probing the role of galaxy interactions in shaping high-energy astrophysical phenomena.
2 Data Analysis
We analyze the aformentioned galaxy merger catalog that contains 328,151 candidate mergers (Ackermann et al., 2018). They classified galaxy mergers in SDSS imaging data using deep convolutional neural networks (CNNs) (Krizhevsky et al., 2017), specifically adopting the Xception architecture (Chollet, 2016). Their approach employed transfer learning by initializing network weights from models pre-trained on the ImageNet dataset (Deng et al., 2009), which is a huge dataset of a few million natural images from thousands of categories, like cats, dogs and many others. The classifier was trained on a labeled dataset derived from SDSS Data Release 7 (Abazajian et al., 2009), consisting of RGB JPEG cutouts obtained via the SDSS image cutout service 222https://skyserver.sdss.org/dr12/en/help/docs/api.aspx. Ground truth labels were sourced from the Galaxy Zoo project (Lintott et al., 2008, 2011), which offers crowd sourced morphological classifications based on visual inspection by citizen scientists. The positive class comprised 3,003 visually confirmed mergers from the catalog of Darg et al. (2010), selected using a weighted merger vote fraction and limited to the redshift range . The negative class included 10,000 randomly selected non interacting galaxies from the Galaxy Zoo project for the same redshift interval, with , serving as a control sample. To mitigate class imbalance during training, stratified sampling was employed to construct balanced mini batches with equal numbers of merger and non-merger galaxies. Model performance was evaluated using 4-fold cross-validation, with class distributions preserved across folds. Training was carried out on subsets of varying sizes (3,000, 1,500, 900, 500, and 300 samples), demonstrating that transfer learning consistently improved performance.
Following training, the CNN was evaluated on a separate test set that was never used during training. For each test galaxy, the model produced a continuous score representing the predicted probability of the galaxy being a merger. A threshold of was used to assign binary labels, galaxies with were classified as mergers, and those with as non-interacting systems. Classification performance was assessed using standard evaluation metrics, including precision, recall, and F1-score. Their method achieved a precision of 0.97, recall of 0.96, and F1-score of 0.97, surpassing the performance of earlier merger classification techniques such as those proposed by earlier works (Goulding et al., 2018; Cotini et al., 2013; Hoyos et al., 2012). The results highlight the high reliability and robustness of the CNN-based approach in identifying galaxy mergers from imaging data. While the CNN classifier demonstrates strong performance in merger identification, several factors may contribute to biases in its classification accuracy. First, the training labels are derived from Galaxy Zoo, a citizen science project involving human visual classification. These labels, while statistically robust in aggregate, are subject to individual perceptual biases and inconsistencies, which may be implicitly learned by the CNN. Second, class definitions are based on thresholds of the merger vote fraction (), with mergers defined by and non-mergers by . Galaxies with intermediate values are excluded, potentially introducing boundary effects and under representing ambiguous cases. Third, transfer learning from the ImageNet domain may introduce domain mismatch. In Figure 1 we showcase the pictorial representation of the workflow followed by Ackermann et al. (2018).
@@ 2 Data Analysis
To isolate high-confidence systems, we applied a strict threshold of across all four classifier outputs, retaining only those galaxies that satisfied this criterion in every fold. This conservative approach yielded a subset of 3,711 robust merger candidates. This stricter criterion minimizes the inclusion of ambiguous cases and emphasizes classifier agreement, albeit at the cost of reducing the final sample size. These selected objects were cross matched with the Sloan Digital Sky Survey Data Release 16 (SDSS DR16) catalog (Ahumada et al., 2020) using TOPCAT (Tool for OPerations on Catalogues And Tables, version 4.10-3), an interactive software tool for manipulating and visualizing large astronomical datasets (Taylor, 2005). TOPCAT facilitates efficient data handling through functionalities like filtering, cross-matching, and plotting, making it ideal for processing astrophysical survey data. A matching radius of 0.1 arcseconds was adopted to ensure precise associations. The resulting cross-match identified 3,185 mergers with SDSS counterparts. Among these matched sources, 1,147 have measured spectroscopic redshifts and 3,166 sources possess SDSS -band apparent magnitudes. We show our analysis workflow in Figure 2.
To characterize the spatial extent of the merger sample, we estimated the total angular sky area over which the 3,166 galaxy mergers are distributed. This effective footprint was computed by dividing the sky into finer bins in right ascension and declination and summing up the area of all the spatial bins that contained at least one merger. The contribution of each bin was corrected for spherical projection by weighting with cos(), where is the central declination of the bin. This produced an estimated sky coverage of approximately , which was used for surface density estimates and association probability calculations. Figure 3 shows the sky distribution in equatorial coordinates using an Aitoff projection.
@@
For the gamma-ray dataset, we utilized 4FGL-DR4 catalog, which contains 7,195 gamma-ray sources detected between 50 MeV and 1 TeV over 14 years of observations (Ballet et al., 2023).
@@
To identify possible associations, we searched for candidate galaxy mergers located within the 4 positional uncertainty ellipse of each 4FGL-DR4 source. This conservative threshold, similar to that used in Zhang et al. (2022), accounts for both Fermi-LAT’s positional uncertainties and uncertainties in the external catalogs. When multiple optical sources were found within a single gamma-ray error ellipse, we assessed the probability of random association using a Poisson-based framework as described below (Browne and Cohen, 1978; Downes et al., 1986; Biggs et al., 2011; Zhang et al., 2022).
@@
The expected number of chance associations ) can be calculated as follow: (Browne and Cohen, 1978; Downes et al., 1986; Biggs et al., 2011; Zhang et al., 2022):
@@
where and are the semi-major and semi-minor axes of the Fermi positional ellipse, and is the surface density of mergers with flux greater than that of the SDSS counterpart having flux equal to . To evaluate , we used the SDSS-derived fluxes computed from the -band magnitudes of the merger candidates. For each candidate to be tested, we determined the number of mergers in the full sample with -band fluxes exceeding the corresponding candidate flux value. The details of the calculation of -band flux from the estimated magnitude can be found in A. Dividing this count by the total sky area covered by the catalog yielded a flux-dependent surface density. This value of was then used in Eq. 1 to calculate the expected number of chance associations (), allowing us to assess the probability of random coincidence.
@@
A small value indicates a low probability of random association, which implies a more reliable counterpart match. We adopt a primary threshold of to define a secure association, and a secondary, more relaxed threshold of to capture additional statistically plausible matches, following the approach of (Zhang et al., 2022). In this work, the authors conducted a systematic cross matching of Fermi-LAT 4FGL sources with multi wavelength catalogs (radio, millimeter, and X-ray) in the South Pole Telescope Sunyaev-Zel’dovich (SPT-SZ) survey region. They adopted a matching radius of 4 positional uncertainty and a -value threshold of to identify statistically plausible associations. They demonstrated that while the probability threshold of ensures high purity, it may exclude real counterparts, especially in crowded fields or when positional uncertainties are large. By extending the threshold to , they achieved a better trade off between completeness and reliability, which maintained low contamination levels under the relaxed criterion.
Given the rarity and low surface density of galaxy mergers, we expect the false match rate to remain low even with a relaxed threshold (Downes et al., 1986; Biggs et al., 2011).
2.1 Data Provenance
The galaxy merger catalog and associated convolutional neural network (CNN) classifications are adopted directly from Ackermann et al. (2018). These include the merger probability scores derived using the Xception architecture and their validation against Galaxy Zoo labels. The contributions of this work consist of: (i) applying a stricter probability threshold () to define a robust subset of mergers, (ii) cross-matching these mergers with SDSS DR16 counterparts, (iii) calculating flux-dependent surface densities and sky area coverage, (iv) statistical validation of merger–gamma-ray associations using a Poisson-based framework, and (v) control-sample analysis to assess chance coincidence rates. This delineation clarifies the division between previously published data products and the new analyses carried out in this study.
3 Results